OpenStudy (anonymous):

find f'(x) f(x)=1/(secx)^7

zepdrix (zepdrix):

\[\Large\bf\sf \sec x\quad=\quad \frac{1}{\cos x}\] Use that^\[\Large\bf\sf \left(\frac{1}{\sec x}\right)^7\quad=\quad ?\]

OpenStudy (anonymous):

\[\frac{ 1 }{ (secx)^7 }\]

zepdrix (zepdrix):

??

zepdrix (zepdrix):

Use the identity.

zepdrix (zepdrix):

\[\Large\bf\sf \sec x\quad=\quad \frac{1}{\cos x}\qquad\implies\qquad \frac{1}{\sec x}\quad=\quad \cos x\]

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{1}{\sec x}\right)^7\quad=\quad ?\]

OpenStudy (anonymous):

so it wont be secx=secxtanx

OpenStudy (anonymous):

cos (x)

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{1}{\sec x}\right)^7\quad=\quad (\cos x)^7\]

OpenStudy (wisp):

Going off of what zepdrix said, you then use the Chain-Chain Rule to find f'(x) = 7(cosx) * - sinx => f'(x) = -7sinxcosx.

OpenStudy (wisp):

*7(cosx)^6 * -sinx => f'(x) = -7(sinx)(cosx)^6